OpenStudy (anonymous):

How much heat (kJ) is needed to raise the temperature of 100.0 grams of water from 25.0ºC to 50.0ºC?

OpenStudy (anonymous):

Hello!

OpenStudy (anonymous):

I already worked the problem, however I seem to be a little confused with what the answer should be. I'll show you the steps.

OpenStudy (anonymous):

q= m * c * delta(T) q= 100 * 4.18 j/g º C * 50ºC - 25ºC q=10450

OpenStudy (anonymous):

q=mct Where q is heat transferred m is the gram of substance c is the specific heat capacity. For water is it 4.18J/g*K t is temperature in Kelvin. Formula for Kelvin is K=C+273.15

OpenStudy (anonymous):

My answer choices are this A) 10450 B) 0.598 C) 1.05 D) 10.5 E) none of the above

OpenStudy (anonymous):

T is temperature difference between the 2 Kelvins temperature

OpenStudy (anonymous):

Yep, you did it correctly!

OpenStudy (anonymous):

My issue is I'm not sure if i have to leave the answer as 10450 or do i have to make it according to the sig figs?

OpenStudy (anonymous):

for example the lowest sigfig is 3 so would i leave it as 10450 or 105?

OpenStudy (anonymous):

It would be 1.05 x 10^4 J You need to use scientific notation

OpenStudy (anonymous):

That makes it 3 sf also.

OpenStudy (anonymous):

Is it require for every answer to use scientific notation?

OpenStudy (anonymous):

No, only if lowest sf is lower than the whole number's sf Like lowest sf is 4 and answer is 102345.23 or 102345 It would be 1.023 x 10^5

OpenStudy (anonymous):

Okay so in my case i got the answer 10450. I need to get it to 3 sf because its the lowest in the problem. I would make it 1.045 but round it to 1.05 making it 1.05 x 10^4?

OpenStudy (anonymous):

To make it 3 sf, you round out the 5 to make it gone.

OpenStudy (anonymous):

Right. So my choices are these A) 10450 B) 0.598 C) 1.05 D) 10.5 E) none of the above

OpenStudy (anonymous):

would the answer be C?

OpenStudy (anonymous):

However there isn't a "x 10^4" or does it make it e?

OpenStudy (anonymous):

You need the scientific notation. So what do you think the answer is now?

OpenStudy (anonymous):

1.05 does not equal 10450 or 10500 but 1.05 x10^4 does.

OpenStudy (anonymous):

I think it's E

OpenStudy (anonymous):

Yes, it would be because the answers are not in correct sf.

OpenStudy (anonymous):

Thanks a lot!